OpenStudy (anonymous):

x8 divided by x4 times x2

OpenStudy (anonymous):

(-6)-2

OpenStudy (australopithecus):

x^(8)/x^(4)*x^(2) = x^(8)/x^(4+2) = x^(8)/x^(6) = x^(8)x^(-6) = x^(8-6) = x^(2)

OpenStudy (australopithecus):

subtract exponents when dividing and add exponents when multiplying similar variables

OpenStudy (australopithecus):

REMEMBER 1/x^(2) = x^(-2)

OpenStudy (anonymous):

 i got this.